Thaddeus and Ian start at the same location and drive in opposite​ directions, but leave at different times. When they are 575 miles​ apart, their combined travel time is 11 hours. If Thaddeus drives at a rate of 50 miles per hour and Ian drives at a rate of 55 miles per​ hour, how long had each been​ driving?
Thaddeus has been driving ____ ​hours, and Ian has been driving ____ hours

Answer by ankor@dixie-net.com(22740)

Thaddeus and Ian start at the same location and drive in opposite​ directions, but leave at different times.
When they are 575 miles​ apart, their combined travel time is 11 hours.
If Thaddeus drives at a rate of 50 miles per hour and Ian drives at a rate of 55 miles per​ hour, how long had each been​ driving?
:
let t = T's travel time
the total of both times is 11 hrs, therefore
(11-t) = I's travel time
:
Write a distance equation, dist = speed * time
T's dist + I's dist = 575 miles
50t + 55(11-t) = 575
50t + 605 - 55t = 575
50t - 55t = 575 - 605
-5t = -30
t = -30/-5
t = +6 hrs is T's travel time
then
11-6 = 5 hrs is I's travel time
:
Check, find actual dist for each
6*50 = 300 mi
5*55 = 275 mi
-------------
total d: 575